# Bounds for the relative class number problem for function fields

### v1

@@ 3.3.1. $q\in\mathopen{}\mathclose{{}\left\{3,4}\right\}$
For $d=2$, we apply Lemma 2.3 and Equations 2.2 and 2.3 to write
$c_{\_}1g+c_{\_}0$$\geq\#C(\mathbb{F}_{\_}q),$
$=q+1-T_{\_}{C,q},$
$=q+1+T_{\_}{A,q},$
@@ 3.3.2. $q=2$
$3.2728\displaystyle 0.8085g+$ 3.2728 $\geq 1.0977(0.7365g+6.5441)-3.9107,$
$\geq 1.0977(a_{\_}1+0.3542\cdot 2a_{\_}2+0.1524\cdot 3a_{\_}3+0.0677\cdot 4a_{\_}4)-3.1329,$
$\geq 1.0977\cdot a_{\_}1+0.2751\cdot 2a_{\_}2-0.1774\cdot 3a_{\_}3+0.0295\cdot 4a_{\_}4-3.1329,$
@@
$\quad+0.1774(2^{3}+1-\#C(\mathbb{F}_{\_}{2^{3}}))+0.0295(\#C(\mathbb{F}_{\_}{2^{4}})-2^{4}-1),$
$=-T_{\_}{C,2}-0.2456\cdot T_{\_}{C,2^{2}}+0.0887\cdot 2T_{\_}{C,2^{3}}-0.0295\cdot T_{\_}{C,2^{4}},$
$=T_{\_}{A,2}+0.2456\cdot T_{\_}{A,2^{2}}+0.0887\cdot T_{\_}{A,2^{3}}+0.0295\cdot T_{\_}{A,2^{4}},$
$\geq 1.1691\dim(A)-1.9169\log\#A(\mathbb{F}_{\_}2),$
@@
$3.2555\displaystyle 0.5993g+$ 3.2555 $\geq 0.8136(0.7365g+6.5441)-2.0688,$
$\geq 0.8136(a_{\_}1+0.3542\cdot a_{\_}2+0.1524\cdot 3a_{\_}3+0.0677\cdot 4a_{\_}4)-2.0688,$
$=0.8136\cdot a_{\_}1-0.2751\cdot 2a_{\_}2+0.0887\cdot 3a_{\_}3-0.0295\cdot 4a_{\_}4-2.0688,$
@@
$\qquad+0.0887(\#C(\mathbb{F}_{\_}{2^{3}})-2^{3}-1)+0.0295(2^{4}+1-\#C(\mathbb{F}_{\_}{2^{4}})),$
$=-T_{\_}{C,2}+0.2456\cdot T_{\_}{C,2^{2}}-0.0887\cdot T_{\_}{C,2^{3}}+0.0295\cdot T_{\_}{C,2^{4}},$
$=T_{\_}{A,2}+0.2456\cdot T_{\_}{A,2^{2}}+0.0887\cdot T_{\_}{A,2^{3}}+0.0295\cdot T_{\_}{A,2^{4}},$
$\geq 1.1691\dim(A)-1.9169\log\#A(\mathbb{F}_{\_}2),$
@@ 5.3. Purely geometric case
In this case, Riemann-Hurwitz gives the inequality
@@ 6.1. Analysis of the Prym
where $E$ is the elliptic curve with #$\#E(\mathbb{F}_{\_}4)=1$ (isogeny class 1.4.ae) and $B$ is an abelian variety not containing $E$ as a factor and satisfying #$\#B(\mathbb{F}_{\_}4)=2$. By [4, Theorem 3.2], we have #$\#B(\mathbb{F}_{\_}4)\geq 2^{\dim(B)}$ and so $\dim(B)=1$; specifically, $B$ is an elliptic curve with #$\#B(\mathbb{F}_{\_}4)=2$ (isogeny class 1.4.ad). We have
@@ 6.3.4. $g\geq 2$ and $d>2$
consequently, $C$ must be the unique curve with #$\#C(\mathbb{F}_{\_}4)=10$ and $J$ belongs to the isogeny class 2.4.f_n. However, in this case #$\#J(\mathbb{F}_{\_}4)=55$ is odd, and so by class field theory (as in [5, (7.5)]) $F$ cannot admit a nonconstant unramified quadratic extension, a contradiction.
@@ 7.3.3. $g\geq 2$ and $d=2$
(Suggested by Everett Howe) Suppose the contrary that the isogeny class contains the Jacobian of a curve $C$. The “resultant 2” criterion [2, Theorem 4.5] shows that $C$ must be a double cover of a curve whose Jacobian belongs to one of the isogeny classes 2.3.d_i or 3.3.d_j_s; however, the former does not contain a Jacobian (e.g., by the “resultant 1” criterion). The latter contains the Jacobian of a unique curve $C_{\_}0$; the double cover $C\to C_{\_}0$ must be étale. We may now run over étale double covers of $C_{\_}0$ to see that none of them have the correct number of $\mathbb{F}_{\_}3$-rational points; this computation can be found in the notebook q3geometric.ipynb listed in Table 4. ∎
@@
  • If #$\#C(\mathbb{F}_{\_}3)=7$, then #$\#C^{\prime}(\mathbb{F}_{\_}3)=0$ and so the covering $C^{\prime}\to C$ cannot have any $\mathbb{F}_{\_}3$-rational ramification points. As in [5, (7.6), (7.7)], we must have $2\#C(\mathbb{F}_{\_}3)\leq\#C^{\prime}(\mathbb{F}_{\_}{3^{2}})=\#C(\mathbb{F}_{\_}{3^{2}})-T_{\_}{A,3^{2}}$ and so #$\#C(\mathbb{F}_{\_}{3^{2}})\geq 2\#C(\mathbb{F}_{\_}3)+T_{\_}{A,3^{2}}\geq 17$. However, this LMFDB query shows that the possible isogeny classes for the Jacobian of a curve $C$ with #$\#C(\mathbb{F}_{\_}3)=7$ are 2.3.df, 2.3.dg, 2.3.dh, and this shows that #$\#C(\mathbb{F}_{\_}{3^{2}})\leq 16$.
@@ 7.3.5. $g\geq 2$ and $d>3$
We first rule out the possibility that $F^{\prime}/F$ admits an intermediate subfield $F^{\prime\prime}$; note that this includes the case where $F^{\prime}/F$ is cyclic. In this case, the extension $F^{\prime\prime}/F$ would have relative class number 1 or 2; however, the former is impossible because [5, Table 4] implies #$\#C(\mathbb{F}_{\_}3)\in\{3,5\}$, while the latter is impossible since Lemma 7.2 implies #$\#C(\mathbb{F}_{\_}3)\leq 6$.
We assume hereafter that $F^{\prime}/F$ has no intermediate subfield, and set notation as in section 2.5; in particular, $F^{\prime\prime}/F$ is now the Galois closure of $F^{\prime}/F$. As in [6, Lemma 4.5], we can use the point counts of $C$ and $C^{\prime}$ to see that the Galois group of $F^{\prime\prime}/F$ must contain a 4-cycle, and so must equal $S_{\_}4$.
If the quadratic resolvent $F^{\prime\prime\prime}/F$ is constant, then $C_{\_}{\mathbb{F}_{\_}9}$ admits a cyclic cubic étale cover which is not a base extension from $C$, so as in [6, Remark 5.2] we have #$\#J(\mathbb{F}_{\_}9)/\#J(\mathbb{F}_{\_}3)\equiv 0\pmod{3}$; this excludes the isogeny classes 2.3.dg, 2.3.dh, 2.3.ei, 2.3.ek. We handle the remaining cases using splitting sequences as in [6].
@@
If the quadratic resolvent is purely geometric, then it corresponds to an étale double cover $C^{\prime\prime\prime}\to C$. Let $E^{\prime}$ be the Prym of this cover; again as in [6, Remark 5.2], the existence of an étale cyclic triple cover of $C^{\prime\prime\prime}$ not induced from $C$ ensures that #$\#E^{\prime}(\mathbb{F}_{\_}3)\equiv 0\pmod{3}$. By Lemma 7.2 we cannot have $T_{\_}{E^{\prime},3}=-2$; this leaves only the option $T_{\_}{E^{\prime},3}=1$. Since $C^{\prime\prime\prime}\to C$ is a Galois cover, #$\#C^{\prime\prime\prime}(\mathbb{F}_{\_}3)=\#C(\mathbb{F}_{\_}3)-T_{\_}{E^{\prime},3}$ must be even, forcing #$\#C(\mathbb{F}_{\_}3)=7$. By class field theory, #$\#J(\mathbb{F}_{\_}3)$ must be even; this restricts $J$ to the single isogeny class 2.3.dg, so
@@ 8. Explicit calculations for geometric covers
This exhaust requires as input a list of systems of polynomial equations defining isomorphism class representatives for the curves of genus $g$ over $\mathbb{F}_{\_}q$. As of this writing, this data is available in LMFDB for all pairs $(q,d)$ thtat we require; however, some notes on sourcing are in order.

### v2

@@ 3.3.1. $q\in\mathopen{}\mathclose{{}\left\{3,4}\right\}$
For $d=2$, we apply Lemma 2.3 and Equations 2.3 and 2.4 to write
$c_{\_}1g+c_{\_}0$$\geq\#C(\mathbb{F}_{\_}q)$
$=q+1-T_{\_}{C,q}$
$=q+1+T_{\_}{A,q}$
@@ 3.3.2. $q=2$
$3.2728\displaystyle 0.8085g+$ 3.2728 $\geq 1.0977(0.7365g+6.5441)-3.9107$
$\geq 1.0977(a_{\_}1+0.3542\cdot 2a_{\_}2+0.1524\cdot 3a_{\_}3+0.0677\cdot 4a_{\_}4)-3.1329$
$\geq 1.0977\cdot a_{\_}1+0.2751\cdot 2a_{\_}2-0.1774\cdot 3a_{\_}3+0.0295\cdot 4a_{\_}4-3.1329$
@@
$\quad+0.1774(2^{3}+1-\#C(\mathbb{F}_{\_}{2^{3}}))+0.0295(\#C(\mathbb{F}_{\_}{2^{4}})-2^{4}-1)$
$=-T_{\_}{C,2}-0.2456\cdot T_{\_}{C,2^{2}}+0.0887\cdot 2T_{\_}{C,2^{3}}-0.0295\cdot T_{\_}{C,2^{4}}$
$=T_{\_}{A,2}+0.2456\cdot T_{\_}{A,2^{2}}+0.0887\cdot T_{\_}{A,2^{3}}+0.0295\cdot T_{\_}{A,2^{4}}$
$\geq 1.1691\dim(A)-1.9169\log\#A(\mathbb{F}_{\_}2)$
@@
$3.2555\displaystyle 0.5993g+$ 3.2555 $\geq 0.8136(0.7365g+6.5441)-2.0688$
$\geq 0.8136(a_{\_}1+0.3542\cdot a_{\_}2+0.1524\cdot 3a_{\_}3+0.0677\cdot 4a_{\_}4)-2.0688$
$=0.8136\cdot a_{\_}1-0.2751\cdot 2a_{\_}2+0.0887\cdot 3a_{\_}3-0.0295\cdot 4a_{\_}4-2.0688$
@@
$\qquad+0.0887(\#C(\mathbb{F}_{\_}{2^{3}})-2^{3}-1)+0.0295(2^{4}+1-\#C(\mathbb{F}_{\_}{2^{4}}))$
$=-T_{\_}{C,2}+0.2456\cdot T_{\_}{C,2^{2}}-0.0887\cdot T_{\_}{C,2^{3}}+0.0295\cdot T_{\_}{C,2^{4}}$
$=T_{\_}{A,2}+0.2456\cdot T_{\_}{A,2^{2}}+0.0887\cdot T_{\_}{A,2^{3}}+0.0295\cdot T_{\_}{A,2^{4}}$
$\geq 1.1691\dim(A)-1.9169\log\#A(\mathbb{F}_{\_}2)$
@@ 5.3. Purely geometric case
In this case, Riemann–Hurwitz gives the inequality
@@ 6.1. Analysis of the Prym
where $E$ is the elliptic curve with #$\#E(\mathbb{F}_{\_}4)=1$ (isogeny class 1.4.ae) and $B$ is an abelian variety not containing $E$ as a factor and satisfying #$\#B(\mathbb{F}_{\_}4)=2$. By [5, Theorem 3.2], we have #$\#B(\mathbb{F}_{\_}4)\geq 2^{\dim(B)}$ and so $\dim(B)=1$; specifically, $B$ is an elliptic curve with #$\#B(\mathbb{F}_{\_}4)=2$ (isogeny class 1.4.ad). We have
@@ 6.3.4. $g\geq 2$ and $d>2$
consequently, $C$ must be the unique curve with #$\#C(\mathbb{F}_{\_}4)=10$ and $J$ belongs to the isogeny class 2.4.f_n. However, in this case #$\#J(\mathbb{F}_{\_}4)=55$ is odd, and so by class field theory (as in [6, (7.5)]) $F$ cannot admit a nonconstant unramified quadratic extension, a contradiction.
@@ 7.3.3. $g\geq 2$ and $d=2$
(Suggested by Everett Howe) Suppose the contrary that the isogeny class contains the Jacobian of a curve $C$. The “resultant 2” criterion [3, Theorem 4.5] shows that $C$ must be a double cover of a curve whose Jacobian belongs to one of the isogeny classes 2.3.d_i or 3.3.d_j_s; however, the former does not contain a Jacobian (e.g., by the “resultant 1” criterion). The latter contains the Jacobian of a unique curve $C_{\_}0$; the double cover $C\to C_{\_}0$ must be étale. We may now run over étale double covers of $C_{\_}0$ to see that none of them have the correct number of $\mathbb{F}_{\_}3$-rational points; this computation can be found in the notebook q3geometric.ipynb listed in Table 4. ∎
@@
  • If #$\#C(\mathbb{F}_{\_}3)=7$, then #$\#C^{\prime}(\mathbb{F}_{\_}3)=0$ and so the covering $C^{\prime}\to C$ cannot have any $\mathbb{F}_{\_}3$-rational ramification points. As in [6, (7.6), (7.7)], we must have $2\#C(\mathbb{F}_{\_}3)\leq\#C^{\prime}(\mathbb{F}_{\_}{3^{2}})=\#C(\mathbb{F}_{\_}{3^{2}})-T_{\_}{A,3^{2}}$ and so #$\#C(\mathbb{F}_{\_}{3^{2}})\geq 2\#C(\mathbb{F}_{\_}3)+T_{\_}{A,3^{2}}\geq 17$. However, this LMFDB query shows that the possible isogeny classes for the Jacobian of a curve $C$ with #$\#C(\mathbb{F}_{\_}3)=7$ are 2.3.df, 2.3.dg, 2.3.dh, and this shows that #$\#C(\mathbb{F}_{\_}{3^{2}})\leq 16$.
@@ 7.3.5. $g\geq 2$ and $d>3$
We first rule out the possibility that $F^{\prime}/F$ admits an intermediate subfield $F^{\prime\prime}$; note that this includes the case where $F^{\prime}/F$ is cyclic. In this case, the extension $F^{\prime\prime}/F$ would have relative class number 1 or 2; however, the former is impossible because [6, Table 4] implies #$\#C(\mathbb{F}_{\_}3)\in\{3,5\}$, while the latter is impossible since Lemma 7.2 implies #$\#C(\mathbb{F}_{\_}3)\leq 6$.
We assume hereafter that $F^{\prime}/F$ has no intermediate subfield, and set notation as in section 2.5; in particular, $F^{\prime\prime}/F$ is now the Galois closure of $F^{\prime}/F$. As in [7, Lemma 4.5], we can use the point counts of $C$ and $C^{\prime}$ to see that the Galois group of $F^{\prime\prime}/F$ must contain a 4-cycle, and so must equal $S_{\_}4$.
If the quadratic resolvent $F^{\prime\prime\prime}/F$ is constant, then $C_{\_}{\mathbb{F}_{\_}9}$ admits a cyclic cubic étale cover which is not a base extension from $C$, so as in [7, Remark 5.2] we have #$\#J(\mathbb{F}_{\_}9)/\#J(\mathbb{F}_{\_}3)\equiv 0\pmod{3}$; this excludes the isogeny classes 2.3.dg, 2.3.dh, 2.3.ei, 2.3.ek. We handle the remaining cases using splitting sequences as in [7].
@@
If the quadratic resolvent is purely geometric, then it corresponds to an étale double cover $C^{\prime\prime\prime}\to C$. Let $E^{\prime}$ be the Prym of this cover; again as in [7, Remark 5.2], the existence of an étale cyclic triple cover of $C^{\prime\prime\prime}$ not induced from $C$ ensures that #$\#E^{\prime}(\mathbb{F}_{\_}3)\equiv 0\pmod{3}$. By Lemma 7.2 we cannot have $T_{\_}{E^{\prime},3}=-2$; this leaves only the option $T_{\_}{E^{\prime},3}=1$. Since $C^{\prime\prime\prime}\to C$ is a Galois cover, #$\#C^{\prime\prime\prime}(\mathbb{F}_{\_}3)=\#C(\mathbb{F}_{\_}3)-T_{\_}{E^{\prime},3}$ must be even, forcing #$\#C(\mathbb{F}_{\_}3)=7$. By class field theory, #$\#J(\mathbb{F}_{\_}3)$ must be even; this restricts $J$ to the single isogeny class 2.3.dg, so
@@ 8. Explicit calculations for geometric covers
This exhaust requires as input a list of systems of polynomial equations defining isomorphism class representatives for the curves of genus $g$ over $\mathbb{F}_{\_}q$. As of this writing, this data is available in LMFDB for all pairs $(q,d)$ that we require; however, some notes on sourcing are in order.